OpenStudy (anonymous):

The product of a number and the number increased by 9 is 36. What is the larger of the two solutions? Write the number only.

OpenStudy (anonymous):

i dont understand

OpenStudy (shamim):

let the number is x

OpenStudy (gir_lover_♥):

^

OpenStudy (anonymous):

wait so the answer is x?????

OpenStudy (shamim):

if the number increased by 9 then it will b x+9

OpenStudy (anonymous):

x+9? thats the answer??

OpenStudy (anonymous):

math confuses me

OpenStudy (shamim):

i think x(x+9)=36

OpenStudy (anonymous):

ill try it

OpenStudy (shamim):

ok

OpenStudy (anonymous):

not correct.

OpenStudy (shamim):

can u tell me the correct 1

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

i need help jaka

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

i can see that

OpenStudy (anonymous):

i hate math

OpenStudy (anonymous):

same

OpenStudy (thomas5267):

Okay. There are two numbers, \(A\) and \(B\). \(B=A+9\) and \(AB=36\).

OpenStudy (anonymous):

Thank you

OpenStudy (thomas5267):

There exist two \(A\) that satisfies the equation. Find the larger of the two.

OpenStudy (thomas5267):

The question is poorly written though.

OpenStudy (anonymous):

@countryrockbabe

OpenStudy (thomas5267):

Basically substitute \(B=A+9\) into \(AB=36\) and solve the quadratic equation in A.

OpenStudy (anonymous):

Yawn

OpenStudy (thomas5267):

I am not suppose to give the answer according to the code of conduct.

OpenStudy (anonymous):

Give Me The Answer

OpenStudy (anonymous):

I love you foseva

OpenStudy (mathstudent55):

@shamim wrote the correct equation above. Here it is again: x(x+9)=36 Solve it for x. You will get two answers. Pick the larger one.

OpenStudy (shamim):

u know 36=3*12

OpenStudy (shamim):

x^2+9x-36=0 x^2+12x-3x-36=0 right?

OpenStudy (mathstudent55):

Correct so far. Now factor by grouping.

OpenStudy (anonymous):

Yawn